# MMInA: Benchmarking Multihop Multimodal Internet Agents

### v1

@@ Abstract
Autonomous embodied agents live on an Internet of multimedia websites. Can they hop around multimodal websites to complete complex user tasks? Existing benchmarks fail to assess them in a realistic, evolving environment for their embodiment across websites. To answer this question, we present MMInA, a multihop and multimodal benchmark to evaluate the embodied agents for compositional Internet tasks, with several appealing properties: 1) Evolving real-world multimodal websites. Our benchmark uniquely operates on evolving real-world websites, ensuring a high degree of realism and applicability to natural user tasks. Our data includes 1,050 human-written tasks covering various domains such as shopping and travel, with each task requiring the agent to autonomously extract multimodal information from web pages as observations; 2) Multihop web browsing. Our dataset features naturally compositional tasks that require information from or actions on multiple websites to solve, to assess long-range reasoning capabilities on web tasks; 3) Holistic evaluation. We propose a novel protocol for evaluating an agent’s progress in completing multihop tasks. We experiment with both standalone (multimodal) language models and heuristic-based web agents. Extensive experiments demonstrate that while long-chain multihop web tasks are easy for humans, they remain challenging for state-of-the-art web agents. We identify that agents are more likely to fail on the early hops when solving tasks of more hops, which results in lower task success rates. To address this issue, we propose a simple memory augmentation approach replaying past action trajectories to reflect. Our method significantly improved both the single-hop and multihop web browsing abilities of agents. See our code and data at https://mmina.cliangyu.com.
[…]
@@ 1 Introduction
Building embodied agents capable of autonomous behaviors navigating in various environments has been a longstanding and intricate challenge in the realm of artificial intelligence research [27, 54, 9, 41]. One common scenario that necessitates automation involves the interaction with digital interfaces [33, 43], with a particular emphasis on the automation of actions performed on rich Internet websites [40, 50, 14]. Real-world web tasks are usually compositional, involving multihop actions across multiple websites. Achieving this goal requires endowing agents with both long-range planning and multimodal reasoning capabilities. This includes understanding high-level instructions derived from user inputs, planning multihop actions within the web browser environment, encompassing HTML content and visual cues from web resources, and making practical predictions based on the observations. While web agents show promise in addressing single-hop tasks that require only one website according to existing benchmarks [21, 24, 25, 53, 17], we observe that they encounter challenge to solve multihop web tasks, which are prevalent in most realistic scenarios that a user requires information from or acts on multiple websites for a high-level task. (Tab. 3). This gap motivates us to establish a multihop web browsing benchmark to assess the usefulness of Internet agents in natural multihop tasks.
@@
Another gap in web agent research is multimodality. Existing benchmarks pose autonomous agent tasks that rely solely on textual inputs and textual information [53, 8, 50]. However, in real-world scenarios, visual information often plays an indispensable role and cannot be disregarded. For instance, consider the task of ”Help me purchase a blue cotton shirt”, where the color attribute, derived from visual information, becomes crucial in fulfilling the user’s request. What’s more, there is a notable lack of current web-based agent benchmarks for emphasis on assessing the capabilities of comprehending and interacting with both textual and visual inputs, where a significant number of them primarily concentrate on tasks that involve text-based interactions.
To address the two issues above, we present MMInA, a novel benchmark designed to advance the field of multihop and multimodal Internet task solving. Our benchmark uniquely operates on evolving real-world websites (Fig. 1), ensuring a high degree of realism and applicability to natural scenarios, (Tab. 2). Central to MMInA is its focus on realistic tasks that users commonly engage in, such as navigating e-commerce platforms, extracting and synthesizing information from content-rich sites like Wikipedia, and performing comparative analysis across multiple web sources. Our carefully designed human-written 1,050 tasks not only challenge the agents to comprehend multimodal inputs at multiple website hops but also to execute sophisticated, multi-step reasoning – a significant leap from the simpler tasks commonly seen in existing benchmarks.
Our extensive experiments with state-of-the-art agents reveal that while there is significant progress in handling simple textual tasks, the integrated and sequential nature of tasks in MMInA poses a substantial challenge. For instance, the best-performing standalone model, GPT-4V [29], achieves an overall success rate of 21.8% across tasks. This is a notable improvement over textual agent baselines but still lags behind the human performance (96.3%). We identify that agents are more likely to fail on the early hops when solving tasks of more hops (Tab. 4), which results in lower task success rates. These results underscore the complexities of real-world web navigation and decision-making, emphasizing the need for further advancements in multimodal and multihop reasoning. To bridge this gap, we propose a simple memory augmentation approach replaying past action trajectories to reflect. Our method significantly improved both the single-hop and multihop web browsing abilities of agents. Memory augmentation is a flexible model-agnostic technique, which can be lent to more large multimodal models as agents in the future.
@@
  • We introduce MMInA, an Internet agent benchmark encompassing 1,050 multihop multimodal tasks spanning 14 diverse websites with evolving and realistic features. The inclusion of multihop tasks enables the assessment of more intricate and human-like actions, closely mirroring real-world problem-solving trajectories. We benchmark the current representative large language models (LLMs) and large multimodal models (LMMs) as agents on the benchmark.
  • We propose a holistic evaluation method for multihop tasks. Based on experimental observations, the task success rates of completing multihop tasks for existing agents are exceedingly low. Our new protocol evaluates the completion results by hops, providing a fine-grained method to assess an agent’s ability to complete multihop tasks.
  • We propose a memory-augmented method that can effectively improve agent performance by reflecting on the action histories of past tasks. The method is simple and flexible, yet effective; significantly improving LMMs as Internet agents.
@@
MiniWoB++ [22] 1 / 1.00 Static simplified websites  (Open-ended) 100
WebShop [49] 1 / 1.00 Static simplified websites  (Open-ended) 1
Mind2Web [8] 1 / 1.00 Static real-world websites  (MC) 131
RUSS [47] 2 / 1.10 Static real-world websites  (MC) 22
WebArena [53] 2 / 1.06 Static real-world websites  (Open-ended) 6
VWA [17] 2 / 1.05 Static real-world websites  (Open-ended) 3
@@ 2 Related Works
Most existing works evaluate autonomous agents on curated textual I/O interfaces, leaving a gap in assessing their performances on real-world automation tasks. APIBench [21], introduced by Gorilla[32], is a tool-augmented LLM benchmark to assess the tool utilization abilities of agents for code generation tasks. AgentBench [24] stepped forward to provide a more general toolkit with lots of closed-box environments to assess agents’ performances in answering user queries. BOLAA [25] is another benchmark that coordinates multiple autonomous agents to make collective decisions. OpenAGI [12] and GAIA [28] are multimodal benchmarks crafted for generalist agents that define multi-step tasks across multiple modalities. However, none of the above benchmarks explored the usage of LLMs or LMMs in web browsing environments or posed an effective evaluation metric specifically tailored for web agent tasks.
@@
WoB [40] and its tool extension MiniWoB++ [22] established a platform of website widgets where agents can complete online tasks through basic keyboard and mouse operations. Webshop [49] was a simulated e-commerce environment featuring 1.18 million real-world products, complemented by 12,087 crowdsourced textual instructions. It converted the action prediction task into a choice-based imitation learning process, which facilitated the accuracy and ability for task execution. However, this approach failed to evaluate open-ended agent actions in the real world. It was also limited by its monotonous design of a one-website environment, which resulted in only a single category of web browsing tasks. Mind2Web [8] tried to construct a generalist web agent, which created a dataset for crafting and benchmarking the web agents by the ability of instruction following. It proposed a two-stage training to convert the action prediction problem into MCQs. SeeAct [52] was a following work that enabled multimodal information for visually understanding rendered webpages and generating more accurate action plans. WebVoyager [13] was capable of capturing screenshots of web pages and then using JavaScript tools to automatically identify interactive elements based on the types of webpage elements. WebArena [53] deployed a standalone set of multicategory websites in an interactive environment. VisualWebArena [17] was a subsequent project that built upon the foundation of WebArena, introducing the reliance on visual cues into the benchmark’s design. The tasks of existing benchmarks are oversimplified whose completions require a single website, which is highly diverged from the natural web browsing tasks and should originally be designed to multihop over a long-horizon setting.
@@ 3 MMInA Benchmark
Websites Task
VWA
[Uncaptioned image] When did programming language that has the largest variance in salary first appear? Answer using the information from the Wikipedia site in the second tab.
MMInA
[Uncaptioned image] Do both LIYFF-Stools Modern Home Office Chair and GGHHJ Adjustable Rotating Salon Stool PU Leather Round Rolling Stool Round Chair have armrest?
Table 2: Comparison between VWA [17] and MMInA. MMInA requires multimodal inputs at multiple steps to accomplish the task, which makes a more challenging multimodal benchmark.
@@ 3.1 Environment
Following [53], we formulate web browsing as a partially observable Markov decision process $\langle S,A,P,R\rangle$. The state space $S$ is the whole Internet content, the status of the browser environment and agent, exceeding representable expressions in practice. Therefore, we pass the partial observation space $\Omega$ of $S$ to the agent. At each time step $t$, an agent arrives at a certain state of a particular web page. The accessibility tree of the screenshot with linked images, with the action/state histories, forms a partial observation $o_{t}\in\Omega$ for the agent. Then the web agent takes an action $a_{t}$ sampled from the action space $A$, being either an executable operation on the web page or a textual output as the answer to a question (Sec. 3.4). The state transition probability matrix $P:S\times A\rightarrow S^{\prime}$ is implicitly encoded as the world knowledge of the Internet environment, which can be inferred or learned by a web agent. Our reward function $R$ is expressed with language output, PASS or FAIL, as a result of each hop, not step. Naturally, we define hop as a subtask that is completed on a specific website. For example, in the task of Fig. 1, the agent receives a PASS if it finds the correct destination in the first hop, another PASS for arriving at the desired flight search page at the second hop, and so on. The task is successful only if the agent receives PASS at each hop.
@@
To achieve a more faithful replication of human actions on the web, we follow [17] to condense the potential agent-executed actions into a set of 12 summarized actions. Leveraging the playwright library, we simulate web pages on an X graphics server, employing a diverse array of actions to interact with the pages. These actions span a broad spectrum of behaviors mirroring human interactions with web pages, encompassing actions such as clicking on links, scrolling up and down using the scroll wheel, typing with the keyboard, and more. From the hop counts shown in Fig. 3, we observe that as the number of hops increases, the number of actions required by the agent also increases. However, it’s worth noting that the average number of actions required for 5-hop data is lower than that for 4-hop data. In our dataset, 4-hop content involves comparative operations, such as ”Which one has a Ferris wheel in the center of the city, Tianjin or Chengdu?” Since our definition of multi-hop tasks involves navigation across different web pages, we do not categorize these comparative questions as 5-hop tasks. On average, an MMInA task takes 12.9 actions to complete.
@@
The observation space $\Omega$ usually embeds partial observations of the Internet to simulate real web browsing experiences. Observations include the task descriptions and the website contents. To represent web content in a reasonable length, we primarily use accessibility trees that provide a structured and unified layout of the web page content. Each node of the accessibility tree has an element ID, the type of the element, and the textual content of the element. If the element is an image, the environment downloads the image and paints the element ID on the image as a reference for the multimodal agent.
@@ 3.2 Multimodal Web Content
Our work at MMInA focuses on multimodality-reliant tasks, which require both images and textual data to complete. For example, the task ”Which one is more furry, Hi&Yeah Comfy Faux Fur Cute Desk Chair or Armen Living Diamond Office Chair?” requires the agent to locate and compare specified items on referenced web pages, analyzing the images and textual web page content to provide an answer.
MMInA’s approach contrasts with VWA, as all tasks in our framework necessitate the processing of both visual and textual information in multiple turns (Tab. 2). With richer multimodal content, MMInA becomes a more challenging yet realistic benchmark.
To tackle these multimodal tasks effectively, MMInA integrates an automated process for extracting the accessibility tree from web pages, encompassing both images and text. This enables the agent to download referenced images and utilize them alongside the accessibility tree as inputs for a multimodal agent. Such an approach underscores the intricate interplay between visual and textual data in solving real-world tasks within a multimodal framework.
@@ 3.3 Multihop Cross-website Browsing
MMInA dataset features multihop tasks across 14 distinct websites, encompassing a diverse range of domains such as shopping, ticket booking, travel guide exploration, and local food discovery (refer to Fig. 2). We define a ”multihop task” as a task that necessitates actions across multiple websites. It requires the agent to automatically go to the next website after finishing a hop. This setting emphasizes the complexity and real-world relevance of multihop web browsing, as they more accurately simulate the sequence of actions a human user typically performs when facing a compositional high-level task. In practice, we provide the links to the available websites (see details in Appendix A) in the task descriptions.
@@ 3.4 Evaluation
Following [53], we have utilized two distinct evaluation methodologies for single-hop tasks, based on the semantics and effectiveness of the predicted actions. Each method is tailored to suit different task characteristics within the MMInA dataset, providing either strict or loose bounds for the evaluation process.
The first method, ”must_include”, is grounded in a keyword-based approach. Here, we establish a set of compositional keywords for each task. The agent’s response is deemed successful (PASS) only if it incorporates all these specified keywords. Conversely, the omission of even a single keyword results in the task being classified as a failure (FAIL). This method ensures a stringent, keyword-focused evaluation of the agent’s performance.
The second method, designated as "fuzzy_match", leverages the advanced capabilities of large language models, exemplified by GPT-3.5-Turbo. This method involves inputting both the reference answer and the agent’s response into the language model. Subsequently, the model is prompted with a query framed as follows: ”Given the statement {pred}, would it be correct to infer {reference}? Yes or No”. Here {pred} represents for agent’s response and {reference} means the reference answer. The evaluation outcome is determined based on the model’s response: a task is adjudged as (PASS) if the model confirms the inference (Yes), and as (FAIL) if it does not (No). This approach offers a more nuanced and flexible evaluation framework, effectively accommodating the subtleties and complexities inherent in language interpretation. To illustrate, in a task where color identification is involved, if the reference answer is ”gold” and the agent’s response is ”yellow”, the "fuzzy_match" method will try to grasp and assess the semantic meaning of the context and provide the conclusive matching result.
@@
In our experiments with multihop tasks(Tab. 3), we often observe a remarkably low completion rate for the entire task, if not zero. To provide a more holistic evaluation, we propose an evaluation method tailored for $N$-hop problems:
The evaluation involves maintaining a queue containing the conditions of each hop’s completion. In particular, the last element of a queue is always an ”END” marker that signifies the whole multihop task is completed, making the queue’s length $N+1$. An agent succeeds at a hop once it finds out the required information (e.g., an answer string) or reaches the desired state (e.g., a specific URL). For simplicity, our benchmark enforces that the agent completes tasks in sequence, i.e., the agent is only allowed to proceed to the next hop if the current hop is correctly completed. A task is completed only if all the hops are correctly completed in sequence.
Our single-hop and multihop evaluation methods aim to provide a systematic and insightful approach to assess the performance of agents in tackling multihop tasks, addressing the challenges posed by such tasks’ complexity.

### v2

@@ Abstract
Autonomous embodied agents live on an Internet of multimedia websites. Can they hop around multimodal websites to complete complex user tasks? Existing benchmarks fail to assess them in a realistic, evolving environment for their embodiment across websites. To answer this question, we present MMInA, a multihop and multimodal benchmark to evaluate the embodied agents for compositional Internet tasks, with several appealing properties: 1) Evolving real-world multimodal websites. Our benchmark uniquely operates on evolving real-world websites, ensuring a high degree of realism and applicability to natural user tasks. Our data includes 1,050 human-written tasks covering various domains such as shopping and travel, with each task requiring the agent to extract multimodal information from web pages as observations autonomously; 2) Multihop web browsing. Our dataset features naturally compositional tasks that require information from or actions on multiple websites to solve, to assess long-range reasoning capabilities on web tasks; 3) Holistic evaluation. We propose a novel protocol for evaluating an agent’s progress in completing multihop tasks. We experiment with both standalone (multimodal) language models and heuristic-based web agents. Extensive experiments demonstrate that while long-chain multihop web tasks are easy for humans, they remain challenging for state-of-the-art web agents. We identify that agents are more likely to fail on the early hops when solving tasks with more hops, which results in lower task success rates. To address this issue, we propose a simple memory augmentation approach that replays past action trajectories to reflect. Our method significantly improves the performance of both the single-hop and multihop web browsing abilities. Our code and data are available on github.com/shulin16/MMInA.
footnotetext: $\ast$ Equal Contribution.     $\dagger$ Project Lead.    🖂 Corresponding Author.
@@ 1 Introduction
Building embodied agents capable of autonomous behaviors navigating in various environments has been a longstanding and intricate challenge in the realm of artificial intelligence research Maes (1993); Ziemke (1998); Florian (2003); Steels and Brooks (2018). One common scenario that necessitates automation involves the interaction with digital interfaces Puig et al. (2018); Toyama et al. (2021), with a particular emphasis on the automation of actions performed on rich Internet websites Shi et al. (2017); Yao et al. (2023); Hong et al. (2024). Real-world web tasks are inherently compositional, requiring multihop actions across multiple websites. To accomplish this, agents must possess both long-range planning and multimodal reasoning capabilities. This includes understanding high-level instructions from user inputs, planning multihop actions across the web browser environment by leveraging HTML content and visual cues, and making informed predictions based on observations. While web agents perform well on single-hop tasks that require interactions with only one website, according to existing benchmarks Li et al. (2023b); Liu et al. (2024, 2023b); Zhou et al. (2024); Koh et al. (2024), we observe that most agents struggle with multihop web tasks, which are prevalent in real-world scenarios where users must gather information from or take actions across multiple websites to accomplish a high-level task (Tab. 2). This gap motivates us to establish a multihop web browsing benchmark to assess the usefulness of Internet agents in natural multihop tasks.
@@
Another gap in web agent research is multimodality. Existing benchmarks pose autonomous agent tasks that rely solely on textual information Zhou et al. (2024); Deng et al. (2023); Yao et al. (2023). However, in real-world scenarios, visual information often plays an indispensable role and cannot be disregarded. For example, consider the task of “Help me purchase a blue cotton shirt”, where the color attribute, derived from visual information, becomes crucial in fulfilling the user’s request. However, there is a notable lack of current web-based agent benchmarks with emphasis on assessing the capabilities of comprehending and interacting with both textual and visual inputs, where a significant number of them primarily concentrate on tasks that involve text-based interactions.
To address the two issues above, we present MMInA, a novel benchmark designed to advance multihop and multimodal Internet task solving. MMInA operates on real-world, evolving websites (Fig. 1), offering high realism and applicability to natural scenarios (Tab. A3). It focuses on realistic tasks users commonly perform, such as navigating e-commerce platforms, extracting information from content-rich sites like Wikipedia, and conducting comparative analysis across web sources. Our 1,050 human-written tasks challenge agents to process multimodal inputs across multiple website hops and execute complex, multi-step reasoning, surpassing the simpler tasks found in existing benchmarks.
Our experiments with state-of-the-art models as agents’ reasoning backbones show that while progress has been made in handling simple textual tasks, the integrated and sequential nature of tasks in MMInA remains a significant challenge. For example, GPT-4V Achiam et al. (2023) achieves a 21.8% success rate, which improves upon textual baselines but falls short of human performance (96.3%). Agents often fail early in multi-hop tasks (Tab. 3), leading to lower success rates. These results highlight the complexity of real-world web navigation and decision-making, underscoring the need for further advancements in multimodal and multihop reasoning. To bridge this gap, we propose a memory augmentation approach that replays past action trajectories, significantly improving both single-hop and multihop performance. This model-agnostic technique can be applied to other large multimodal models in the future.
@@
  • We introduce MMInA, an Internet agent benchmark featuring 1,050 multihop multimodal tasks across 14 diverse, evolving websites with realistic features. The multihop tasks assess more complex, human-like problem-solving actions, closely resembling real-world scenarios. We evaluate current large language models (LLMs) and large multimodal models (LMMs) as agents’ backbones on the benchmark.
  • We propose a holistic evaluation method for multihop tasks. Building on the limitations of task success rate evaluation, which often yields poor agent performance and limited insights, we propose a new protocol for multihop tasks based on hop success rate. This provides a more granular assessment, offering deeper insights into the relationship between agent behavior and hop length, facilitating a more informed analysis of long-chain reasoning capabilities.
  • We propose a lightweight and adaptable memory-augmented method that enhances agent performance by leveraging past action histories.
@@
MiniWoB++ Liu et al. (2018) 1 / 1.00 Static simplified websites  (Open-ended) 100
WebShop Yao et al. (2022) 1 / 1.00 Static simplified websites  (Open-ended) 1
Mind2Web Deng et al. (2023) 1 / 1.00 Static real-world websites  (MC) 131
RUSS Xu et al. (2021) 2 / 1.10 Static real-world websites  (MC) 22
WebArena Zhou et al. (2024) 2 / 1.06 Static real-world websites  (Open-ended) 6
VWA Koh et al. (2024) 2 / 1.05 Static real-world websites  (Open-ended) 3
WebVoyager He et al. (2024) 4 / 2.40 Dynamic real-world websites  (Open-ended) 15
@@ 2 Related Works
Most existing works evaluate autonomous agents on curated textual I/O interfaces, leaving a gap in assessing their performance on real-world automation tasks. Tool-use benchmarks like Li et al. (2023b); Patil et al. (2024); Liu et al. (2024) aim to assess agents’ performances with tool-usage capability; BOLAA Liu et al. (2023b) is another benchmark that coordinates multiple autonomous agents to make collective decisions. OpenAGI Ge et al. (2023) and GAIA Mialon et al. (2024) are multimodal benchmarks crafted for generalist agents that define multi-step tasks across multiple modalities. However, none of the above benchmarks explored the usage of LLMs or LMMs in web browsing environments or posed an effective evaluation metric specifically tailored for web agent tasks.
@@
Webshop Yao et al. (2022) builds a simulated e-commerce environment featuring 1.18 million real-world products, complemented by 12,087 crowdsourced textual instructions, while Mind2Web Deng et al. (2023), CogAgent Hong et al. (2024), and SeeAct Zheng et al. (2024) try to construct a generalist web agent; WebVoyager He et al. (2024) can automatically identify interactive elements based on the types of webpage. Recently, WebArena Zhou et al. (2024) deploys a standalone set of multicategory websites in an interactive environment, where VisualWebArena Koh et al. (2024) is a subsequent project that built upon WebArena, introducing the reliance on visual cues into the benchmark’s design. However, we found that the tasks of existing benchmarks are oversimplified whose completions requiring a single website, which is highly diverged from the natural web browsing tasks and should originally be designed for multihop over a long-horizon setting.
@@ 3.1 Environment
Following Zhou et al. (2024), we formulate web browsing as a partially observable Markov decision process $\langle S,A,P,R\rangle$. The state space $S$ is the whole Internet content, the status of the browser environment and agent, exceeding representable expressions in practice. Therefore, we pass the partial observation space $\Omega$ of $S$ to the agent. At each time step $t$, an agent arrives at a certain state of a particular web page. The accessibility tree of the screenshot with linked images, with the action/state histories, forms a partial observation $o_{t}\in\Omega$ for the agent. Then the web agent takes an action $a_{t}$ sampled from the action space $A$, being either an executable operation on the web page or a textual output as the answer to a question (Sec. 3.5). The state transition probability matrix $P:S\times A\rightarrow S^{\prime}$ is implicitly encoded as the world knowledge of the Internet environment, which can be inferred or learned by a web agent. Our reward function $R$ is expressed with language output, PASS or FAIL, as a result of each hop. Naturally, we define one hop as a subtask that is completed on a specific website. For example, in the task of Fig. 1, the agent receives a PASS if it finds the correct destination in the first hop, another PASS for arriving at the desired flight search page at the second hop, and so on.
@@
We follow Koh et al. (2024) to condense the potential agent-executed actions into a set of 12 summarized actions. Leveraging the playwright library, we simulate web pages on an X graphics server, employing a diverse array of actions to interact with the pages. These actions span a broad spectrum of behaviors mirroring human interactions with web pages, encompassing actions such as clicking on links, scrolling up and down using the scroll wheel, typing with the keyboard, etc. A higher hop count corresponds to a higher number of actions. On average, an MMInA task takes 12.9 actions to complete.
@@
The observation space $\Omega$ usually embeds partial observations of the Internet to simulate real web browsing experiences. Observations include the task descriptions and the website content. To represent web content in a reasonable length, we primarily use accessibility trees that provide a structured and unified layout of the web page content. Each node of the accessibility tree has an element ID, the type of the element, and the textual content of the element. If the element is an image, the environment downloads the image and paints the element ID on the image as a reference for the multimodal agent.
3.2 Dataset Construction
Data Structure
MMInA adapted question styles from the WebQA dataset and used GPT-4V to generate similar questions with multimodal content (e.g., “help me buy a yellow jacket online”). The prompt structure includes: 1) Instructions: some basic concepts about tasks, accessibility tree, and actions; 2) Rules: such as “do nothing after action [stop]”; 3) QA examples: some tiny examples are here to help the agent to understand instructions and above; 4) Reference Websites: this is the universe of all potential websites the agent may visit; 5) Task: a multihop multimodal task to solve. The hops vary from $1$ to $10$. Additionally, we manually crafted questions to diversify style, scope, and content across categories like shopping, search, and booking. In summary, MMInA tasks combine $2,989$ hops from $14$ dynamic, real-world websites to ensure a diverse and realistic set of challenges.
Annotators & Annotation Protocols
Annotators are human experts in web browsing and vary in age and gender to ensure the fairness of labeling. The annotators first proposed task templates varying in intent and difficulty. Each template generates $2-10$ tasks. All annotators then followed a “minimalist” approach, where annotators, acting as “omniscient readers”, completed tasks using the shortest paths with all crucial website nodes recorded. This annotation protocol enhances the trajectory diversity in the evaluation process, where the ground-truth trajectory should be a subset of any successfully visited trajectories of the agents.
Performance Metrics
Evaluating real websites is challenging due to their dynamic nature, as web content frequently changes. To address this, we propose a new evaluation metric for multihop tasks based on the visited URLs. A task is considered successfully completed only when all required websites are visited in order, ensuring alignment between the agent’s actions and the task objectives. Details are shown in Sec. 3.5.
@@ 3.3 Multimodal Web Content
Our work at MMInA focuses on multimodality-reliant tasks, which require both images and textual data to complete. For example, the task “Which one is more furry, Hi&Yeah Comfy Faux Fur Cute Desk Chair or Armen Living Diamond Office Chair?” requires the agent to locate and compare specified items on referenced web pages, analyzing the images and textual web page content to provide an answer. MMInA’s approach contrasts with VWA, as all tasks in our framework necessitate the processing of both visual and textual information in multiple turns (Tab. A3).
MMInA includes an automated process for extracting accessibility trees from web pages while identifying and downloading the image contents in the current view. This allows agents to use images alongside the accessibility tree as inputs, highlighting the critical interaction between visual and textual data in solving real-world tasks within a multimodal framework.
@@ 3.4 Multihop Cross-website Browsing
MMInA dataset features multihop tasks across 14 distinct websites, covering diverse domains such as shopping, ticket booking, travel guides, and local food discovery (Fig. 2). A “multihop task” requires actions across multiple websites, with the agent automatically moving to the next site after completing each hop. This setup mirrors the complexity and real-world relevance of multihop web browsing, simulating the sequence of actions a human user would typically perform when tackling a high-level task. Task descriptions include links to the available websites (see Appendix B for details).
@@ 3.5 Evaluation
Following Zhou et al. (2024), we implemented two evaluation methodologies for single-hop tasks to assess the semantics and effectiveness of predicted actions within the MMInA dataset. These methods offer either stringent or lenient criteria, depending on the task’s characteristics.
The first method, “must_include”, adopts a keyword-based evaluation approach. For each task, a set of essential keywords is defined. An agent’s response is considered successful (PASS) if it incorporates all these keywords. Missing any keyword results in a failure (FAIL), ensuring a rigorous assessment based on keyword inclusion. The second method, “fuzzy_match”, utilizes the capabilities of advanced language models, such as GPT-3.5-Turbo, to evaluate responses. It compares the agent’s response and a reference answer by asking the model: “Given the statement pred, would it be correct to infer reference? Yes or No”, where pred is the agent’s response and reference is the reference answer. The evaluation is determined by the model’s reply: a PASS is recorded if the answer is “Yes”, and a FAIL if “No”. This method offers a flexible evaluation framework that accommodates linguistic nuances, such as assessing semantic similarities between “gold” and “yellow” in color identification tasks.
@@
In our experiments with multihop tasks (Tab. 2), we often observe a remarkably low success rate for the entire task, if not zero. To provide a more granular evaluation, we propose an evaluation method tailored for $N$-hop problems: The evaluation involves maintaining a queue containing the conditions of each hop’s completion. In particular, the last element of a queue is always an “END” marker that signifies the whole multihop task is completed, making the queue’s length $N+1$. An agent succeeds at a hop once it finds out the required information (e.g., an answer string) or reaches the desired state (e.g., a specific URL). For simplicity, our benchmark enforces that the agent completes tasks in sequence, i.e., the agent is only allowed to proceed to the next hop if the current hop is correctly completed. A task is completed only if all the hops are correctly completed in sequence.
Our interleaved single-hop and multihop evaluation methods aim to provide a systematic and insightful approach to assess the performance of agents in tackling multihop tasks, addressing the challenges posed by such tasks’ complexity.